# Quantum Electrodynamics of graphene Landau levels in a deep-sub-wavelength hyperbolic phonon polariton cavity

### v1

@@ II.1 Theory of the Green’s function of HPP cavities
In free space, the Poisson equation in Eq. (1) lacks a time-dependent component, making $\phi(\bm{r})$ a completely classical field. However, in HPP cavities, the non-trivial frequency dependence of the dielectric permittivity tensor causes the emergence of resonant modesGubbin_JChemPhys_2022 , which can be described as quantum mechanical modes. Consequently, the poles of Eq. (LABEL:eqn:green_function_hBN_expansion_Main) correspond to the quantum mechanical modes of the cavity Hamiltonian, ^$\hat{\mathcal{H}}_{\rm cav}=\sum_{\bm{q}_{\parallel},n_{z},s}\hbar\omega_{{q}_{\parallel},n_{z},s}\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}\hat{a}_{\bm{q}_{\parallel},n_{z},s}$, where the operator $\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$ ($\hat{a}_{\bm{q}_{\parallel},n_{z},s}$) creates (destroys) a cavity excitation. Since the cavity is a deeply sub-wavelength one, the electromagnetic field is well described by a scalar potential
@@
where $S$ is the surface area of the cavity, and $\phi_{{q}_{\parallel},n_{z},s}(z)\equiv\cos(q_{z,n_{z}}z)\sqrt{{2}/({N_{{q}_{\parallel},n_{z},s}L_{z}})}$, with $N_{{q}_{\parallel},n_{z},s}$ being a normalization factor. The procedure to determine this normalization factor will be detailed momentarily. The physical electric field can be obtained by taking the gradient of the potential in Eq. (3). Further details are reported in Subsection III.1. To determine this normalization factor, we employ the following strategy. First, we calculate the quantum scalar Green’s function $g_{\rm q}^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$ using the scalar potential $\hat{\phi}(\bm{r}_{\parallel},z)$ in Eq. (3) and the standard Kubo formulaGiuliani_and_Vignale as detailed in Subsection III.1. As we will show momentarily, the scalar potential $\hat{\phi}(\bm{r}_{\parallel},z)$ expanded in terms of the HPP modes $\hat{a}_{\bm{q}_{\parallel},n_{z},s},\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$ in Eq. (3) does not describe the full electrostatic contribution. Instead, there is a frequency-independent contribution that is not captured by these modes. As only resonant processes contribute to dissipation, all dissipative processes are accounted by $g_{\rm q}^{(0)}({q}_{\parallel},z,z^{\prime},\omega)$, the imaginary part of the quantum Green’s function matches that of the Green’s function derived from the Poisson equation as given in Eq. (LABEL:eqn:green_function_hBN_expansion_Main) i.e. ${\rm Im}~{}g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)={\rm Im}~{}g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$. By imposing this condition we determine the normalization factor $N_{{q}_{\parallel},n_{z},s}$ by imposing that ${\rm Im}~{}g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)={\rm Im}~{}g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$. Still, while the imaginary parts coincide, the real parts of $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ and $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ differ.
Specifically, $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ vanishes in the high-frequency limit, whereas $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})\equiv\lim_{\omega\to\infty}g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ remains finite. As a result, the quantum Green’s function $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$, associated with the modes $\hat{a}_{\bm{q}_{\parallel},n_{z},s}$ and $\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$, satisfies the Kramers-Kronig relations. In contrast, the full scalar Green’s function $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ is not analytic on the complex plane, as $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$ is nonzero. Consequently, $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ fulfills the Kramers-Kronig relations only when $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$ is subtracted. Indeed, the following decomposition holds: $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)=g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)+g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$. The term $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ effectively captures the resonant processes tied to the excitation of the system’s quantum mechanical modes $\hat{a}_{\bm{q}_{\parallel},n_{z},s},\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$, whereas $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$ accounts for a frequency-independent renormalization of the Coulomb interaction. The details of this section are provided in Sec. III.
@@
II.2 Theory of Landau levels in Graphene and their coupling to a subwavelength cavity
@@
We now consider single-layer graphene, which will be later positioned at the center of the HPP cavity (i.e. at $z=0$), subject to a strong, uniform, and static perpendicular magnetic field $B$. In this regime, electronic excitations are inter-Landau-level transitions Goerbig11 ; ando1982 ; roldan2009 ; roldan2010 . In the Landau gauge, the energy spectrum of electrons roaming in this crystal are massless Dirac fermion Landau levels (LLs) given by $\varepsilon_{\lambda,n}=\lambda(\hbar v_{\rm F}/\ell_{B})\sqrt{2n}$, where $\ell_{B}=\sqrt{\hbar c/(eB)}$ is the magnetic length, $e$ is the electron charge, $v_{\rm F}$ is the Fermi velocity in graphene, $k_{x}$ is the electronic wavevector along the $x$-axis, which accounts for the degeneracy of the spectrum, $n$ is a non-negative integer representing the LL index, and $\lambda=\pm$ denotes conduction/valence band states. We denote by the symbol $N_{\rm F}$ the index of the last occupied LL. Here, we concentrate on the charge neutrality point of graphene, choosing $N_{\rm F}=0$. In Fig. 2(b), we plot the non-interacting spectrum of LLs, $-{\rm Im}\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)$, where $\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)$ is the bare density-density response function calculated using the Kubo formula roldan2009 ; Giuliani_and_Vignale . This quantity is fully non-local, valid for all wavevectors ${q}_{\parallel}$, which is particularly relevant as the deep sub-wavelength nature of hBN enables exploration of high wave-vector regions. Specifically, accounting for spatial non-local effects is necessary in the limit of rapidly varying matter fields ($q\sim 0.1~{}{\rm nm}^{-1}$), which has been investigated in SNOM experimentsdai2014 , and also in the presence of a magnetic fieldwehmeier2024 .
@@
The LLs spectrum shows electron-hole pair excitations associated with transitions $\lambda,k_{x},n\to\lambda^{\prime},k_{x}^{\prime},n^{\prime}$ and corresponding transition energies $\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n}$. Due to the $k_{x}$ degeneracy, bare LL transitions do not depend on the momentum transfer ${q}_{\parallel}=k_{x}^{\prime}-k_{x}$ and thus appear as flat lines in the ${q}_{\parallel}-\omega$ plane. In Fig. 2(b), we plot transitions between valence-band states $\lambda=-,n$ and conduction-band states $\lambda^{\prime}=+,n^{\prime}$. Dipole-allowed transitions $n^{\prime}=n\pm 1$ are drawn as solid white lines, while other transitions are shown as white dashed lines. In the numerics, we choose a strong magnetic field in the order of $B\approx 5~{}{\rm T}$ (corresponding to $\ell_{B}\approx 11.5~{}{\rm nm}$). Due to the deep sub-wavelength nature of our cavities, we focus on high wavevectors $q_{\parallel}\approx 0.1~{}{\rm nm}^{-1}$, where all transitions are relevant, as $q_{\parallel}\ell_{B}\sim 1$. The discrete nature of LLs leads to a series of transition frequencies that can couple with the electromagnetic modes of the HPP cavity.
@@
We now move one to discuss the coupling between matter degrees of freedom and cavity modes. In the Coulomb gauge, the LLs and the cavity modes interact through the electrostatic interaction Hamiltonian andolina2024 ; Giuliani_and_Vignale :
@@
where $\hat{\phi}(\bm{r}_{\parallel},z=0)$ is the scalar potential of the cavity reported in Eq. (3) and evaluated at the vertical location of graphene and $\hat{n}(\bm{r}_{\parallel})$ is the electron density operator in graphene. This Hamiltonian describes the coupling of the electronic density in graphene with the electric field of the cavity, leading to the formation of hybrid light-matter states known as polaritons hopfield1958 ; Kavokin11 . However, the interaction Hamiltonian in Eq. (4) only accounts for the dynamical part of the field, corresponding to the quantum Green’s function $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$. Therefore, it is necessary to include the effect of the statically-screened potential, $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$. This can be incorporated into the Hamiltonian in the form of a standard two-body screened Coulomb interaction, ^$\hat{\mathcal{H}}_{\rm C}=({1}/{2S})\sum_{\bm{q}_{\parallel}}\hat{n}_{-\bm{q}_{\parallel}}e^{2}g_{\infty}^{(0)}(\bm{q}_{\parallel},0,0)\hat{n}_{\bm{q}_{\parallel}}$, where $\hat{n}_{\bm{q}_{\parallel}}$ is the momentum-resolved density operator. In the previous summation, the ${\bm{q}}_{\parallel}={\bm{0}}$ contribution is canceled by the inclusion of a positive background (Jellium) to ensure charge neutrality Giuliani_and_Vignale . Since our analysis focuses on finite $q_{\parallel}$, this procedure does not affect the results.

### v2

@@ II.1 Theory of the Green’s function of HPP cavities
In free space, the Poisson equation in Eq. (1) lacks a time-dependent component, making $\phi(\bm{r})$ a completely classical field. However, in HPP cavities, the non-trivial frequency dependence of the dielectric permittivity tensor causes the emergence of resonant modesGubbin_JChemPhys_2022 , which can be described as quantum mechanical modes. Consequently, the poles of Eq. (LABEL:eqn:green_function_hBN_expansion_Main) correspond to the quantum mechanical modes of the cavity Hamiltonian, ^$\hat{\mathcal{H}}_{\rm cav}=\sum_{\bm{q}_{\parallel},n_{z},s}\hbar\omega_{{q}_{\parallel},n_{z},s}\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z}}\hat{a}_{\bm{q}_{\parallel},n_{z},s}$, where the operator $\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$ ($\hat{a}_{\bm{q}_{\parallel},n_{z},s}$) creates (destroys) a cavity excitation. Since the cavity is a deeply sub-wavelength one, the electromagnetic field is well described by a scalar potential
@@
where $S$ is the surface area of the cavity, and $\phi_{{q}_{\parallel},n_{z},s}(z)\equiv\cos(q_{z,n_{z}}z)\sqrt{{2}/({N_{{q}_{\parallel},n_{z},s}L_{z}})}$ are the amplitude coefficients of the modes, with $N_{{q}_{\parallel},n_{z},s}$ being a normalization factor. The procedure to determine this normalization factor will be discussed momentarily and detailed in Subsection III.1. Eq. (3) is one of our key results, as it enables the solution of the Poisson equation (Eq. (1)) to be expressed in terms of creation and annihilation operators. This formulation allows for a direct quantification of quantum vacuum fluctuations of the field. The corresponding physical electric field associated with the scalar potential is obtained by taking the gradient of Eq. (3). The normalization factors are determined by ensuring that the imaginary part of the quantum scalar Green’s function $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$, calculated using the scalar potential $\hat{\phi}(\bm{r}_{\parallel},z)$ in Eq. (3) and the standard Kubo formulaGiuliani_and_Vignale , matches that of the classical Green’s function $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$, which is derived from the Poisson equation. However, their real parts differ: $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ vanishes at high frequencies, while $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})\equiv\lim_{\omega\to\infty}g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ remains finite. Consequently, $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ satisfies the Kramers-Kronig relations, whereas $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)$ does so only after subtracting $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$, thus $g^{(0)}(q_{\parallel},z,z^{\prime},\omega)=g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)+g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$. This decomposition separates resonant quantum mechanical contributions $g_{\rm q}^{(0)}$, associated with the system’s excitation modes $\hat{a}_{\bm{q}_{\parallel},n_{z},s},\hat{a}^{\dagger}_{\bm{q}_{\parallel},n_{z},s}$, from a static Coulomb interaction renormalization encoded in $g_{\infty}^{(0)}$. All details of this procedure are provided in Sec. III.
@@
II.2 Theory of Landau levels in Graphene and their coupling to a sub-wavelength cavity
@@
We now consider single-layer graphene, which will be later positioned at the center of the HPP cavity (i.e. at $z=0$), subject to a strong, uniform, and static perpendicular magnetic field $B$. In this regime, electronic excitations are inter-Landau-level transitions Goerbig11 ; ando1982 ; roldan2009 ; roldan2010 . In the Landau gauge, the energy spectrum of electrons roaming in this crystal are massless Dirac fermion Landau levels (LLs) given by $\varepsilon_{\lambda,n}=\lambda(\hbar v_{\rm F}/\ell_{B})\sqrt{2n}$, where $\ell_{B}=\sqrt{\hbar c/(eB)}$ is the magnetic length, $e$ is the electron charge, $v_{\rm F}$ is the Fermi velocity in graphene, $k_{x}$ is the electronic wavevector along the $x$-axis, which accounts for the degeneracy of the spectrum, $n$ is a non-negative integer representing the LL index, and $\lambda=\pm$ denotes conduction/valence band states. We denote by the symbol $N_{\rm F}$ the index of the last occupied LL. Here, we focus on the charge neutrality point of graphene, choosing $N_{\rm F}=0$. In Fig. 2(b), we plot the non-interacting spectrum of LLs, $-{\rm Im}\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)$, where $\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)$ is the bare density-density response function calculated using the Kubo formula roldan2009 ; Giuliani_and_Vignale .
This quantity is fully non-local, valid for all wavevectors ${q}_{\parallel}$, which is particularly relevant as the deep sub-wavelength nature of hBN enables exploration of high wave-vector regions. Specifically, accounting for spatial non-local effects is necessary in the limit of rapidly varying matter fields ($q\sim 0.1~{}{\rm nm}^{-1}$), which has been investigated in SNOM experimentsdai2014 , and also in the presence of a magnetic fieldwehmeier2024 .
The LLs spectrum shows electron-hole pair excitations associated with transitions $\lambda,k_{x},n\to\lambda^{\prime},k_{x}^{\prime},n^{\prime}$ and corresponding transition energies $\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n}$. Due to the $k_{x}$ degeneracy, bare LL transitions do not depend on the momentum transfer ${q}_{\parallel}=k_{x}^{\prime}-k_{x}$ and thus appear as flat lines in the ${q}_{\parallel}-\omega$ plane. The magnitude of the bare density-density current response function $\chi^{(0)}_{\rm nn}({q}_{\parallel},\omega)$ critically depends on the matrix elements of the electronic density operator
$\begin{split}\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})\equiv\sum_{k^{\prime}_{x},k_{x}}\lvert\braket{k_{x}^{\prime},n^{\prime},\lambda^{\prime}}{\hat{n}_{\bm{q}_{\parallel}}}{k_{x},n,\lambda}\rvert^{2}~{},\\ \end{split}$ (4)
@@
where $|k_{x},n,\lambda\rangle$ are the electronic state associated with the quantum number $\lambda,k_{x},n$ and $\hat{n}_{\bm{q}_{\parallel}}$ is the momentum resolved electronic density. In Fig. 2(b) the transitions between valence-band states $\lambda=-,n$ and conduction-band states $\lambda^{\prime}=+,n^{\prime}$ are shown. Dipole-allowed transitions $n^{\prime}=n\pm 1$ are drawn as solid white lines, while other transitions are shown as white dashed lines. In the numerics, we choose a strong magnetic field in the order of $B\approx 5~{}{\rm T}$ (corresponding to $\ell_{B}\approx 11.5~{}{\rm nm}$). Due to the deep sub-wavelength nature of our cavities, we focus on high wavevectors $q_{\parallel}\approx 0.1~{}{\rm nm}^{-1}$, where all transitions are relevant, as $q_{\parallel}\ell_{B}\sim 1$. The discrete nature of LLs leads to a series of transition frequencies that can couple with the electromagnetic modes of the HPP cavity.
@@
We now move on to discuss the coupling between matter degrees of freedom and cavity modes. In the Coulomb gauge, the LLs and the cavity modes interact through the electrostatic interaction Hamiltonian andolina2024 ; Giuliani_and_Vignale :
@@
where $\hat{\phi}(\bm{r}_{\parallel},z=0)$ is the scalar potential of the cavity reported in Eq. (3) and evaluated at the vertical location of graphene and $\hat{n}(\bm{r}_{\parallel})$ is the electron density operator in graphene. This Hamiltonian describes the coupling of the electronic density in graphene with the electric field of the cavity, leading to the formation of hybrid light-matter states known as polaritons hopfield1958 ; Kavokin11 . However, the interaction Hamiltonian in Eq. (5) only accounts for the dynamical part of the field, corresponding to the quantum Green’s function $g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)$. Therefore, it is necessary to include the effect of the potential, $g_{\infty}^{(0)}(q_{\parallel},z,z^{\prime})$. This can be incorporated into the Hamiltonian in the form of a standard two-body screened Coulomb interaction,
^$\hat{\mathcal{H}}_{\rm C}=\frac{1}{2S}\sum_{\bm{q}_{\parallel}}\hat{n}_{-\bm{q}_{\parallel}}e^{2}g_{\infty}^{(0)}(\bm{q}_{\parallel},0,0)\hat{n}_{\bm{q}_{\parallel}}~{},$ (6)
where $\hat{n}_{\bm{q}_{\parallel}}$ is the momentum-resolved density operator. In the previous summation, the ${\bm{q}}_{\parallel}={\bm{0}}$ contribution is canceled by the inclusion of a positive background (Jellium) to ensure charge neutrality Giuliani_and_Vignale . Since our analysis focuses on finite $q_{\parallel}$, this procedure does not affect the results. Eq. (6) represents the Coulomb interaction unaffected by the phononic resonances of hBN, $\omega_{i}^{\rm T},\omega_{i}^{\rm L}$. However, the term $g_{\infty}^{(0)}(\bm{q}_{\parallel},0,0)$ accounts for the screening effects from higher-frequency modes in hBN, encapsulated in $\epsilon_{i}(\infty)$, as well as the influence of the metallic mirrors, which is incorporated through the boundary conditions of Eq. (1).
@@
II.3 The polaritonic spectrum and the vacuum Rabi frequency
[…]
Figure 4: (Color online) Panels (a,b) display the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}=(\Omega_{q_{\parallel},n_{z},s,-,0,+,1}/\omega_{q_{\parallel},n_{z},s})$, where $\Omega_{q_{\parallel},n_{z},s,-,0,+,1}$ denotes the Rabi frequency for the Landau level (LL) transition $(0,+1)$, and $\omega_{q_{\parallel},n_{z},s}$ is the cavity mode frequency. This ratio is shown as a function of $L_{z}$ (in ${\rm nm}$) for different $n_{z}$ for a fixed $\bar{q}_{\parallel}=0.1{\rm nm^{-1}}$ and $B=10{\rm T}$. Panels (a) show these quantities for the upper Reststrahlen band using blue dotted curves, while panels (b) present the lower Reststrahlen band with red dashed curves. The line thickness diminishes as the quantum number $n_{z}$ increases.
Alternatively, one can diagonalize the total Hamiltonian ^$\hat{\cal{H}}_{\rm tot}=\hat{\cal{H}}_{\rm LLs}+\hat{\cal{H}}_{\rm C}+\hat{\cal{H}}_{\rm cav}+\hat{\cal{H}}_{\rm int}$. In the single excitation manifold, electron-hole excitations between LLs can be treated as bosonic quasiparticles ciuti2005 . In this limit, the total Hamiltonian reduces to a Hopfield Hamiltonianhopfield1958 , a quadratic bosonic Hamiltonian that describes light-matter excitations and can be exactly diagonalized in terms of polaritonic modes. The polariton energies obtained in this procedure are identical to those that one finds from the zeros of dynamical dielectric function $\epsilon_{\rm RPA}({q}_{\parallel},\omega)$.
Fig. 3 shows the electronic spectrum at the neutrality point $N_{\rm F}=0$. Panel (a) presents this spectrum as a function of the wavevector $q_{\parallel}$, while panel (b) displays the spectrum as a function of $(\hbar v_{\rm F}/\ell_{B})$. The latter spectrum can be experimentally obtained by varying the external magnetic field $B$ zhang2005 . Inside the Reststrahlen bands (defined by $\omega^{\rm T}_{i}<\omega<\omega^{\rm L}_{i}$, with these frequencies indicated by black dashed lines), the LL transitions hybridize with HPP cavity modes. The vacuum Rabi frequencies of the system represent the interaction strength between a selected LL transition and a cavity mode. As discussed below, one of these couplings can reach up to $10\%$ of the cavity frequency, placing our system in the ultra-strong coupling regimeKockum19 ; geiser2012 . Furthermore, a single LL transition hybridizes with multiple cavity modes, a regime known as super-strong coupling kuzmin2019 , leading to a complex polaritonic structure beyond simple two-transition models. This region of the spectrum is mainly controlled by the LL transitions encoded in ${\chi}^{(0)}_{\rm nn}({q}_{\parallel},\omega)$ and by the quantum Green’s function associated with the cavity modes ${g}^{(0)}_{\rm q}({q}_{\parallel},0,0,\omega)$, corresponding to the expansion in Eq. (3). Outside the Reststrahlen bands, LL transitions experience a $q_{\parallel}$-dependent positive frequency shift, known as the depolarization shift Giuliani_and_Vignale ; shtrichman2001 , leading to finite dispersion as apparent in Fig. 3(a). Panel (b) shows that, even off-resonance, the LL transitions do not coincide with the non-interacting case (indicated by white solid and dashed lines). This shift originates from the frequency-independent Green’s function $g_{\infty}^{(0)}({q}_{\parallel},0,0)$ rather than from the HPP poles, as it occurs outside the Reststrahlen bands, where $\omega\neq\omega_{{q}_{\parallel},n_{z},s}$ This effect is largely due to the confinement induced by the two metallic mirrors. Indeed, this effect is significantly suppressed in the limit $L_{z}\to\infty$ (see Fig. 9(c) in Sect. V and the related discussion), which corresponds to removing the metallic mirrors and filling the entire space with hBN. Interestingly, despite originating from the Coulomb interactions, this positive shift is reminiscent of the effect of the diamagnetic term associated with transverse modes, which induces a blue shift in polaritonic frequencies.
As detailed in Section V, our theory allows to compute the vacuum Rabi frequency $\Omega_{q_{\parallel},{n}_{z},{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}}$, defined as the coupling between the cavity mode $q_{\parallel},{n}_{z},{s}$ and the inter-LLs transition $n,\lambda\to n^{\prime},\lambda^{\prime}$ given by:
$\hbar\Omega_{q_{\parallel},{n}_{z},{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}}=\phi_{{q}_{\parallel},n_{z},s}(0)\sqrt{{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}~{},$ (8)
where $\phi_{{q}_{\parallel},n_{z},s}$ are the mode’s amplitude coefficients defined by Eq. (3) while ${{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}$ are the matrix elements of the electronic density operator in Eq. (4). In standard C-QED dalibard1982 ; DDB_arxiv_2023 ; ciuti2005 , it is well known that the Rabi frequency is proportional to the product of the vacuum fluctuations of the electric field $\sqrt{\langle\hat{E}^{2}\rangle}$ and the dipole moment of the atom in the cavity, $d$, such that $\Omega_{\rm R}\sim d\sqrt{\langle\hat{E}^{2}\rangle}$. Eq. (8) is the most important result of this article as it generalizes this concept to non-local ($q_{\parallel}\neq 0$), sub-wavelength cavities interacting with a 2D material. Indeed, this equation shows how the coupling depends on the vacuum fluctuations of the cavity field and of the electronic density. In Fig. 4, we present the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}\equiv(\Omega_{\bar{q}_{\parallel},n_{z},s,-,0,+,1}/\omega_{\bar{q}_{\parallel},n_{z},s})$ as a function of $L_{z}$ (in ${\rm nm}$) for different $n_{z}$ while $\bar{q}_{\parallel}=0.1{nm}^{-1}$. Notably, this quantity exhibits a non-monotonic dependence on $L_{z}$, reaching a maximum for a specific $L_{z}$ while vanishing in both limits $L_{z}\to 0$ and $L_{z}\to\infty$. This behavior stands in stark contrast to the naive expectation that the Rabi frequency should scale as $1/\sqrt{L_{z}}$, an assumption that underlies the phenomenological use of the compression factor gao2020 ; schlawin2019 ; wang2019 ; nguyen2023 ; zezyulin2022 ; lenk2022 ; arwas2023 ; pavosevic2022 . While this scaling holds in the regime $\bar{q}_{\parallel}L_{z}\gg 1$, it fails for small $L_{z}$, where the coupling instead vanishes. As apparent from Fig. 4, we observe that the maximum ratio $R_{\bar{q}_{\parallel},n_{z},s,-,0,+,1}$, achieved for $n_{z}=0$, exceeds $10\%$ for the upper Reststrahlen band and approaches $10\%$ for the lower band. This ratio places the system firmly within the ultra-strong coupling regime, where the coupling strength is a significant fraction of the cavity frequency Kockum19 . Notably, as further discussed in Section V, the ultra-strong coupling regime is achieved at a finite wavevector $\bar{q}_{\parallel}$, emphasizing the necessity of employing a non-local theory. This result underscores the limitations of local approximations and highlights the critical role of spatial dispersion in accurately describing light-matter interactions in sub-wavelength cavities. Additionally, a single LL hybridizes with multiple cavity modes, as the Rabi frequencies of many different modes remain non-negligible. This situation characterizes a regime known as super-strong coupling kuzmin2019 .
We now examine the spectrum for finite doping. Fig. 5 shows the electronic spectrum for a doped system with $N_{\rm F}=5$, corresponding to a filling factor $\nu=22$, as $\nu=4N_{\rm F}+2$. At finite charge density $N_{\rm F}\neq 0$ and strong magnetic fields, a plasmonic excitation appears in the system, usually referred to as magnetoplasmon. Magnetoplasmons Giuliani_and_Vignale ; roldan2009 ; muravev2016 arise in doped semiconductors under a magnetic field due to the coupling between plasma oscillations of free carriers and their cyclotron resonance. In Fig. 5(a) we show the spectrum without the contribution of the quantum mechanical modes (which is obtained by setting $g_{\rm q}^{(0)}({q}_{\parallel},0,0,\omega)=0$), revealing the expected magnetoplasmon dispersionroldan2009 . Fig. 5(b) is obtained by taking into account the full cavity contribution. Here, the magnetoplasmon strongly hybridizes with the cavity modes achieving a Rabi spitting of $\approx 25\%$ of the cavity frequency. This significant coupling arises from the well-known fact that most of the oscillator strength, which quantifies the ability of a system to absorb or emit electromagnetic radiation, is associated with the magnetoplasmon resonanceGiuliani_and_Vignale . Finally, we note that in the $L_{z}\to\infty$ limit the Rabi splitting shown in Fig. 5(c) remains of a similar order. Although the coupling between LLs decreases with reduced compression, the corresponding increase in the density of cavity modes compensates, resulting in a splittingDai2015 ; Kumar2015 comparable to the case in which $L_{z}$ is finite. This effect is peculiar of sub-wavelength cavities, as transverse contributions from the vector potential $\hat{\bm{A}}$ vanish in the limit $L_{z}\to\infty$ ashida2023 . Similar considerations hold for the cavity spectrum, $S_{\rm cav}({q}_{\parallel},\omega)$. The details of this section are provided in Sect. V.